What is the magnetic flux for a loop with an area of 0.2m2 is in a uniform magnetic field of 0.5T if the plane containing the loop is tilted at an angle of 60 degrees with respect to the magnetic field vectors? What is the magnetic flux for angles of 0 and 90 degrees.
Answer #1

a] We know that magnetic flux phi = B*area * sin theta where theta is angle between plane and field.

= 0.5*0.2*sin 60 degree

= 0.0866 Wb

b] when theta is zero degree, magnetic flux phi = B*area * sin theta = 0.5*0.2*sin 0 = 0 Wb

c] when theta is 90 degree, magnetic flux phi = B*area * sin theta = 0.5*0.2*sin 90 degree = 0.1 Wb